# Helfrich cylinder Alexander Meiners, Hannes Uecker

### v1

@@ 2.2 Well-posedness
Given an initial condition $u_{0}\in{\cal X}\subset L^{2}({\cal C}_{L}(r))$, i.e., $u_{0}=u_{0}(x,\varphi)$ with periodic boundary conditions in $x$, where ${\cal X}=H^{m}_{\text{per}}({\cal C}_{L}(r))$ with the choice $m\in(3,4)$ explained below, we search for solutions $(t,x,\varphi))\mapsto(u(t,x,\varphi),\Lambda(t))$ of (15) on $[0,T)\times{\cal C}_{r}(L)$. Here ${\cal C}_{L}(r)$ is topologically equivalent to a torus, and the periodic Sobolev spaces of order $m\geq 0$ are
@@
where $\hat{v}_{k}=\left\langle v,e^{ik\cdot(x,\varphi)}\right\rangle$. Clearly, $H^{m_{1}}_{\text{per}}\subset H^{m_{2}}_{\text{per}}$ for $m_{1}\geq m_{2}$, and $[H^{m_{1}}_{\text{per}},H^{m_{2}}_{\text{per}}]_{\theta}=H^{(1-\theta)m_{1}+\theta m_{2}}_{\text{per}}$ for the associated real interpolation spaces, see [LM72] and [tr78, §2.4.2], and $\nabla\in{\cal L}(H^{m}_{\text{per}},H^{m-1}_{\text{per}})$.
@@
with some $0<\varepsilon<r$. Then ${\cal C}_{L}(r)-uN$ is an embedding of ${\cal C}_{L}(r)$, and we can easily state local Lipschitz properties of the operators involved in (15). For convenience we recall some Sobolev embedding results, which hold for compact manifolds with boundary as for compact domains in ${\mathbb{R}}^{n}$, see, e.g., [Aubin82]. [^1] The situation of non compact manifolds is more complicated, but if the manifold is complete, has a positive injection radius and bounded sectorial curvature, then the same results holds as well. In fact, in the following analysis near ${\cal C}_{L}(r)$ we do not strictly need the manifold versions of Sobolev embeddings due to the global coordinates (10) also underlying (22), and hence we basically state the more general embedding results for possible generalization to other manifolds. In 2D, $H^{m}(X)\hookrightarrow L^{\infty}(X)$ for $m>1$ embeds continuously, while $H^{1}(X)$ does not. Also, for $m=1+\alpha$, $H^{m}(X)\hookrightarrow C^{0,\alpha}(X)$, hence the point evaluation $u(x,\phi)$ and in particular the sets $\{|u(x,\varphi)|<r-\varepsilon\}$ are well defined for $m>1$.
@@ 2.2.1 Well-posedness of the PDE part
cf.[LS16], where our $B(u)$ appears as the principal part of the 4th order operator $P(u)$ in [LS16]. In particular, the symbol $\sigma[B(u)](x,\phi,\xi)$ satisfies
and hence $B(u)$ is uniformly elliptic (in the 4th order sense) for $\|u\|_{\infty}<r-\varepsilon$, i.e. there exists constants $c_{1},c_{2}$ such that $c\leq\sigma[B(u)](x,\phi,\xi)\leq C$ for all $(x,\phi,\xi)\in{\cal C}_{L}\times{\mathbb{R}}^{2}$ with $\|\xi\|=1$.
We want to apply [MRS13, Theorem 3.1], which treats abstract quasilinear parabolic problems of the form
@@
[…]
@@
By [MRS13, Theorem 3.1], (24) is then locally well posed in $V$, i.e., there exists a $t_{0}>0$ and a unique solution $u\in C([0,t_{0}),V)\cap H^{1}((0,t_{0}),X_{0})\cap L^{2}((0,t_{0}),H^{4})$. To show (W1) we use Sobolev embeddings and the structures of $B$ and $C$, while (W2) essentially follows from (29).
[…]
@@
To show (W2) we use standard theory for sectorial operators [Henry81, Lun95]. Let $u\in V^{m}_{\varepsilon}$. Since $D(B(u))=H^{4}$ is dense in $X=L^{2}$, it is sufficient to show that $B(u)$ is sectorial in $L^{2}$, i.e., that the spectrum of $B(u)$ is contained in a sector $S_{\alpha,\omega}$ in ${\mathbb{C}}$ with tip at some $\omega\in{\mathbb{R}}$ of angle $\alpha<\pi$ opening to the left, and that there exists and $M>0$ such that the resolvent estimate
@@
holds for all $\lambda\not\in S_{\alpha,\omega}$. Essentially, this follows from the uniform ellipticity (29). In more detail, we show that $B(u)$ induces a coercive sesquilinear form on $H^{2}$, from which (W2) follows via the Lax–Milgram Lemma, see, e.g., [SE24, Corollary 2.29].
[…]
@@
[…]
@@
with $c_{1}>0$ from (29). Proceeding similarly for the other terms from (28) and choosing suitable $\delta_{1,2}$ and then $\omega$ sufficiently large yields (32), uniformly in $u\in V^{m}_{\varepsilon}$. ∎
In summary we have verified (W1) and (W2) with $V=V^{m}_{\varepsilon}$ with $m\in(3,4)$ corresponding to $\vartheta\in(3/4,1)$ for the interpolation space $[L^{2},H^{4}]_{\vartheta,2}=H^{m}$, and we can apply [MRS13, Theorem 3.1] to obtain local existence for (24).
Theorem 2.1.
@@ 2.2.2 The constrained flows
Following Remark 2.1, in order to take the constraints into account, we first consider only one active constraint namely $q_{1}(u)=0$, and fix $\lambda_{2}$, while the same arguments also apply for the case of active $q_{2}$ with fixed $\lambda_{1}$. Defining $\overline{H}=\frac{H}{\|H\|}_{L^{2}}$ there exists a $\sigma_{1}$ such that $-(-{\partial}_{u}E+\lambda_{2}{\partial}_{u}{\cal V}+\lambda_{1}H)=-(-{\partial}_{u}E+\lambda_{2}{\partial}_{u}{\cal V}+\sigma_{1}\overline{H})$, and we define the projection ${\cal P}_{1}:L^{2}({\cal C}_{r}(L))\rightarrow\mathrm{span}\{{\partial}_{u}{\cal A}(u)\}^{\bot}$ by ${\cal P}_{1}f=f-\left\langle\overline{H},f\right\rangle\overline{H}$, and consider
@@
$\lambda_{1}$ $=\frac{1}{\int_{{\cal C}_{r}(L)}H^{2}\,{\rm d}S}\left\langle{\partial}_{u}E{-}\lambda_{2}{\partial}_{u}{\cal V},2H\right\rangle=\frac{1}{\int_{{\cal C}_{r}(L)}H^{2}\,{\rm d}S}\int_{{\cal C}_{r}(L)}(\Delta H{+}2H(H^{2}{-}K){+}c_{0}K{-}2c_{0}H{-}\lambda_{2}{\partial}_{u}{\cal V})H\,{\rm d}S$
@@
Importantly, the term $|\nabla H|^{2}$ includes third order derivatives of $u$ but only quadratically in $\nabla H$ with lower order “coefficients” and hence $\int|\nabla H(u)|^{2}\,{\rm d}S$ is well defined and analytic for $u\in H^{m}_{\text{per}}$ for $m>3$. This yields the analogous result as Theorem 2.1.
@@ 3.1 Linearization and eigenvalues
We consider the linearization of (7) around ${\cal C}_{L}(r)$. Since $G(\cdot,\Lambda):V^{4}_{\varepsilon}\to L^{2}$ is an analytic map, we use (11), (12) and (13) to calculate the Frechét derivative via the Gateaux differential in direction $v$, i.e., ${\partial}_{u}G(0,\Lambda)v=\frac{\,{\rm d}}{\,{\rm d}\varepsilon}G(\varepsilon v,\Lambda)|_{\varepsilon=0}$.
[…]
@@
First recall that $\Delta_{0}={\partial}_{x}^{2}+\frac{1}{r^{2}}{\partial}_{\varphi}^{2}$, the Laplace-Beltrami operator on $\mathcal{C}_{r}$. Using (11) with $\varepsilon h$ and sorting by $\varepsilon$ powers gives
$H(\varepsilon h)=-\frac{\varepsilon(r^{3}h_{xx}+rh_{\varphi\varphi})+O(\varepsilon^{2})}{(r^{2}+\varepsilon 2rh+O(\varepsilon^{2}))^{3/2})}-\frac{1}{(r^{2}+\varepsilon 2rh+O(\varepsilon^{2}))^{1/2}}.$
@@
$\frac{\,{\rm d}}{\,{\rm d}\varepsilon}H(\varepsilon h)=-\frac{(r^{3}h_{xx}+rh_{\varphi\varphi})(r^{2}+O(\varepsilon))^{3/2}+O(\varepsilon)}{(r^{2}+\varepsilon 2rh+O(\varepsilon^{2}))^{3}}-\frac{rh}{(r^{2}+O(\varepsilon))^{3/2}}.$
@@
$\frac{\,{\rm d}}{\,{\rm d}\varepsilon}K(\varepsilon h)=\frac{\,{\rm d}}{\,{\rm d}\varepsilon}\frac{-\varepsilon r^{3}h_{xx}+O(\varepsilon^{2})}{(r^{2}+2\varepsilon 2rh+O(\varepsilon^{2}))^{2}}=\frac{-r^{3}h_{xx}((r^{2}+\varepsilon 2rh+O(\varepsilon^{2}))^{2}+O(\varepsilon)}{(r^{2}+\varepsilon 2rh+O(\varepsilon^{2}))^{4}}.$
@@
$\frac{\,{\rm d}}{\,{\rm d}\varepsilon}\Delta_{\varepsilon h}H(\varepsilon h)|_{\varepsilon=0}=\left(\frac{\,{\rm d}}{\,{\rm d}\varepsilon}\Delta_{\varepsilon h}H(0)\right)\bigg{|}_{\varepsilon=0}+\Delta_{0}\left(\frac{\,{\rm d}}{\,{\rm d}\varepsilon}H(\varepsilon h)\bigg{|}_{\varepsilon=0}\right)=\Delta_{0}\left(\frac{\,{\rm d}}{\,{\rm d}\varepsilon}H(\varepsilon h)\bigg{|}_{\varepsilon=0}\right),$
since $H(0)$ is a constant and $\Delta_{\varepsilon h}$ is non degenerate for small perturbations. ∎
Inserting the expressions from subsection 3.1 we obtain
[…]
@@
Over the unbounded ${\cal C}(r)$, ${\partial}_{u}G(0,\Lambda))$ has a continuous spectrum, but by restricting to the compact cylinder ${\cal C}_{L}(r)$ of length $L$, ${\partial}_{u}G(0,\Lambda))$ has a discrete spectrum with a simple basis of eigenfunctions.
[…]
The eigenvalue equation
@@
$\psi_{mn}(x,\varphi)=e^{{\rm i}(k_{m}x+n\varphi)}$, $k_{m}=2m\pi/L$, with (39a)
$\mu(m,n,\Lambda)=-\frac{1}{2}\Big{[}k_{m}^{2}+\frac{n^{2}}{r^{2}}\Big{]}^{2}+\left(r\lambda_{2}+\frac{1}{r^{2}}\right)\Big{[}k_{m}^{2}+\frac{n^{2}}{r^{2}}\Big{]}$ $-\left(\frac{\lambda_{2}}{r}+\frac{1}{2r^{4}}\right)-\left(\frac{1}{r}-2c_{0}\right)\frac{k_{m}^{2}}{r}.$ (39b)
@@
we obtain $\mu=\frac{1}{2}\Big{[}k_{m}^{2}+\frac{1}{r^{2}}n^{2}\Big{]}^{2}+\left(r\lambda_{2}+\frac{1}{r^{2}}\right)\Big{[}k_{m}^{2}+\frac{1}{r^{2}}n^{2}\Big{]}-\left(\frac{\lambda_{2}}{r}+\frac{1}{2r^{4}}\right)-\frac{1}{r}\left(\frac{1}{r}-2c_{0}\right)r^{-1}k_{m}^{2}$, hence (39) after some rearranging. Clearly $\mu(0,\pm 1,\Lambda)=0$, hence $\sin(\varphi),\cos(\varphi)\in\ker({\partial}_{u}G(0,\Lambda))$. ∎
@@
The domain of ${\partial}_{U}{\cal G}(0,\Lambda)$ is $H^{4}_{\text{per}}({\cal C}_{L}(r))\times{\mathbb{R}}^{2}$, equipped with the norm $\|U\|_{{\partial}_{(u,\Lambda)}\mathcal{G}}=\|u\|_{H^{4}}+\|\Sigma\|_{{\mathbb{R}}^{2}}$, and altogether we have the generalized eigenvalue problem $\Psi=(\psi,\sigma_{1},\sigma_{2})$
${\partial}_{U}\mathcal{G}(0,\Lambda)\Psi=\mu\begin{pmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix}\Psi,$ (41)
with eigenvalues $\mu\in\{\nu(m,n),\,(m,n)\in{\mathbb{Z}}^{2}\backslash\{(0,0)\}\cup\{0\}$ and eigenfunctions
$\Psi_{m,n}=\begin{pmatrix}\psi_{m,n}\\ 1\\ -2\end{pmatrix},\,(m,n)\in{\mathbb{Z}}^{2}\backslash\{(0,0)\},\text{ and }\Psi_{0,0}=\begin{pmatrix}0\\ 1\\ -2\end{pmatrix}.$
@@
(a) The eigenvalues $\mu(0,\pm 1,\lambda)=0$ correspond to translations of ${\cal C}_{L}(r)$ in $y$ and $z$ directions. Each eigenvalue of the form $\mu(m,0,\Lambda)$ and $\mu(0,n,\Lambda)$ has multiplicity two and every “double” eigenvalue $\mu(m,n,\Lambda)$ with $m\neq 0\neq n$ has multiplicity four with eigenfunctions $\psi_{\pm m\pm n}$.
(b) The eigenvalues $\mu(m,n,\Lambda)$ do not explicitly depend on $\lambda_{1}$, but implicitly. Recall (2), which shows that the parameters are not independent at the cylinder.

### v2

@@ 2.2 Well-posedness
Given an initial condition $u_{0}\in{\cal X}\subset L^{2}({\cal C}_{L}(r))$, i.e., $u_{0}=u_{0}(x,\varphi)$ with periodic boundary conditions in $x$, where ${\cal X}=H^{m}_{\text{per}}({\cal C}_{L}(r))$ with the choice $m\in(3,4)$ explained below, we search for solutions $(t,x,\varphi))\mapsto(u(t,x,\varphi),\Lambda(t))$ of (15) on $0,T)\times{\cal C}_{r}(L)$. Here ${\cal C}_{L}(r)$ is topologically equivalent to a torus, and the periodic Sobolev spaces of order $m\geq 0$ are
@@
where $\hat{v}_{k}=\left\langle v,e^{ik\cdot(x,\varphi)}\right\rangle$. Clearly, $H^{m_{1}}_{\text{per}}{\subset}H^{m_{2}}_{\text{per}}$ for $m_{1}{\geq}m_{2}$, and $H^{m_{1}}_{\text{per}},H^{m_{2}}_{\text{per}}]_{\theta}{=}H^{(1-\theta)m_{1}+\theta m_{2}}_{\text{per}}$ for the associated real interpolation spaces, see [LM72] and [Tri78, §2.4.2], and $\nabla\in{\cal L}(H^{m}_{\text{per}},H^{m-1}_{\text{per}})$.
@@
with some $0<\varepsilon<r$. Then ${\cal C}_{L}(r)-uN$ is an embedding of ${\cal C}_{L}(r)$, and we can easily state local Lipschitz properties of the operators involved in (15). For convenience we recall some Sobolev embedding results, which hold for compact manifolds with boundary as for compact domains in ${\mathbb{R}}^{n}$, see, e.g., [Aub82]. [^4] The situation of non compact manifolds is more complicated, but if the manifold is complete, has a positive injection radius and bounded sectorial curvature, then the same results holds as well. In fact, in the following analysis near ${\cal C}_{L}(r)$ we do not strictly need the manifold versions of Sobolev embeddings due to the global coordinates (10) also underlying (21), and hence we basically state the more general embedding results for possible generalization to other manifolds. In 2D, $H^{m}(X)\hookrightarrow L^{\infty}(X)$ for $m>1$ embeds continuously, while $H^{1}(X)$ does not. Also, for $m=1+\alpha$, $H^{m}(X)\hookrightarrow C^{0,\alpha}(X)$, hence the point evaluation $u(x,\phi)$ and in particular the sets $\{|u(x,\varphi)|<r-\varepsilon\}$ are well defined for $m>1$.
@@ 2.2.1 Well-posedness of the PDE part
cf.[LS16], where our $B(u)$ appears as the principal part of the 4th order operator $P(u)$ in [LS16].
Theorem 2.3.
To prove Theorem 2.3 we apply [MRS13, Theorem 3.1], which treats abstract quasilinear parabolic problems of the form
@@
with Gelfand triple $X_{0}\hookrightarrow{\cal X}\hookrightarrow X_{1}=D(A(u))$, i.e., phase space $\cal X$}. For this we could rewrite (23) as ${\partial}_{t}u{=}\widetilde{B}(u)u{+}\widetilde{C}(u)$ with $\widetilde{B}(u){=}\eta(u)^{-1}B(u)$ and $\widetilde{C}(u){=}\eta(u)^{-1}C(u)$, but since $\eta(u)=(1+u)g^{-1/2}$ contains at most first derivatives it will be clear from the following that we can as well keep $\eta(u)$ in (23) on the lhs and consider the rhs of (23) in the setting of (28). As already indicated above, we choose the base space $X_{0}=L^{2}$ and hence $X_{1}=D(B(u))=H^{4}$, and we need to find an open subset $V=V^{m}_{\varepsilon}$ of the phase space $\cal X$} with $X_{0}\subset{\cal X}\subset X_{1}$ such that
@@
By [MRS13, Theorem 3.1], (23) is then locally well posed in $V$, i.e., there exists a $t_{0}>0$ and a unique solution $u\in C([0,t_{0}),V)\cap H^{1}((0,t_{0}),X_{0})\cap L^{2}((0,t_{0}),H^{4})$. To show (W1) we use Sobolev embeddings and the structures of $B$ and $C$, while (W2) essentially follows from $B(u)$ being uniformly elliptic (in the 4th order sense). In detail, the symbol $\sigma[B(u)](x,\phi,\xi)$ satisfies
and hence for $\|u\|_{\infty}<r-\varepsilon$, there exists constants $c_{1},c_{2}$ such that $c_{1}\leq\sigma[B(u)](x,\phi,\xi)\leq c_{2}$ for all $(x,\phi,\xi)\in{\cal C}_{L}\times{\mathbb{R}}^{2}$ with $\|\xi\|=1$.
Lemma 2.4.
@@
To show (W2) we use standard theory for sectorial operators [Hen81, Lun95]. Let $u\in V^{m}_{\varepsilon}$. Since $D(B(u))=H^{4}$ is dense in $X=L^{2}$, it is sufficient to show that $B(u)$ is sectorial in $L^{2}$, i.e., that the spectrum of $B(u)$ is contained in a sector $S_{\alpha,\omega}$ in $\mathbb{C}$} with tip at some $\omega\in{\mathbb{R}}$ of angle $\alpha<\pi$ opening to the left, and that there exists and $M>0$ such that the resolvent estimate
@@
holds for all $\lambda\not\in S_{\alpha,\omega}$. Essentially, this follows from the uniform ellipticity (29). In more detail, we show that $B(u)$ induces a coercive sesquilinear form on $H^{2}$, from which (W2) follows via the Lax–Milgram Lemma, see, e.g., [Sch24, Corollary 2.29].
Lemma 2.5.
@@
$b_{4,0}(u){\partial}_{x}^{4}w={\partial}_{x}^{2}(b_{4,0}(u){\partial}_{x}^{2}w)-{\partial}_{x}(\widetilde{b}_{4,0}(u){\partial}_{x}^{2}w),$
where $\widetilde{b}_{4,0}(u)={\partial}_{x}b_{4,0}(u)$ is still bounded in $L^{\infty}$, i.e., $\|\widetilde{b}_{4,0}(u)\|_{\infty}\leq C_{4,0}$. Therefore, using $|w_{xx}w_{x}|\leq\frac{\delta_{1}}{2}w_{xx}^{2}+\frac{1}{2\delta_{1}}w_{x}^{2}$, $\delta_{1}>0$, and similarly standard $L^{2}$–interpolation estimates such as $\int_{{\cal C}_{L}(r)}w_{x}^{2}\,{\rm d}S\leq\int_{{\cal C}_{L}(r)}\frac{\delta_{2}}{2}w_{xx}^{2}+\frac{1}{2\delta_{2}}w^{2}\,{\rm d}S$, we obtain, e.g.,
$\left\langle(\lambda-b_{4,0}(u))w,w\right\rangle$ $=\lambda\int w^{2}\,{\rm d}S+\int b_{4,0}(u)w_{xx}^{2}+\widetilde{b}_{4,0}(u)w_{xx}w_{x}\,{\rm d}S$
@@
with $c_{1}>0$ from after (29). Proceeding similarly for the other terms from (27) and choosing suitable $\delta_{1,2}$ and then $\omega$ sufficiently large yields (31), uniformly in $u\in V^{m}_{\varepsilon}$. ∎
In summary we have verified (W1) and (W2) with $V=V^{m}_{\varepsilon}$ with $m\in(3,4)$ corresponding to $\vartheta\in(3/4,1)$ for the interpolation space $L^{2},H^{4}]_{\vartheta,2}=H^{m}$, and Theorem 2.3 now folows from [MRS13, Theorem 3.1].
@@ 2.2.2 The constrained flows
In order to take the constraints into account, we first consider only one active constraint namely $q_{1}(u)=0$, and fix $\lambda_{2}$, while the same arguments also apply for the case of active $q_{2}$ with fixed $\lambda_{1}$. Defining $\overline{H}=\frac{H}{\|H\|}_{L^{2}}$ there exists a $\sigma_{1}$ such that $(-{\partial}_{u}E+\lambda_{2}{\partial}_{u}{\cal V}+\lambda_{1}H)=-(-{\partial}_{u}E+\lambda_{2}{\partial}_{u}{\cal V}+\sigma_{1}\overline{H})$, and we define the projection ${\cal P}_{1}:L^{2}({\cal C}_{r}(L))\rightarrow\mathrm{span}\{{\partial}_{u}{\cal A}(u)\}^{\bot}$ by ${\cal P}_{1}f=f-\left\langle\overline{H},f\right\rangle\overline{H}$, and consider
@@
$\lambda_{1}$ $=\frac{1}{\int_{{\cal C}_{r}(L)}H^{2}\,{\rm d}S}\left\langle{\partial}_{u}E{-}\lambda_{2}{\partial}_{u}{\cal V},2H\right\rangle$
$=\frac{1}{\int_{{\cal C}_{r}(L)}H^{2}\,{\rm d}S}\int_{{\cal C}_{r}(L)}(\Delta H{+}2H(H^{2}{-}K){+}c_{0}K{-}2c_{0}H{-}\lambda_{2}{\partial}_{u}{\cal V})H\,{\rm d}S$
@@
Importantly, the term $|\nabla H|^{2}$ includes third order derivatives of $u$ but only quadratically with lower order “coefficients” and hence $\int|\nabla H(u)|^{2}\,{\rm d}S$ is well defined and analytic for $u\in H^{m}_{\text{per}}$ for $m>3$. This yields the analogous result as Theorem 2.3.
@@ 3 Stability of ${\cal C}_{L}$
We consider the linearization of (5) around ${\cal C}_{L}(r)$. This requires some lengthy computations, but since ${\cal C}_{L}(r)$ has constant mean and Gaussian curvatures (with $K=0$), we can explicitly compute the dispersion relation (Lemma 3.1) and subsequently the stability ranges of and bifurcation points from ${\cal C}_{L}(r)$ (Corollary 3.7).
@@ 3.1 Linearization and eigenvalues
Since $G(\cdot,\Lambda):V^{4}_{\varepsilon}\to L^{2}$ is an analytic map, we use (11), (12) and (13) to calculate the Frechét derivative via the Gateaux differential in direction $v$, i.e., ${\partial}_{u}G(0,\Lambda)v=\frac{\,{\rm d}}{\,{\rm d}\varepsilon}G(\varepsilon v,\Lambda)|_{\varepsilon=0}$; recall that the Laplace-Beltrami operator on ${\cal C}_{L}((r)$ is $\Delta_{0}={\partial}_{x}^{2}+\frac{1}{r^{2}}{\partial}_{\varphi}^{2}$ in our coordinate system.
Lemma 3.1.
@@
Using (11) with $\varepsilon v$ and sorting by $\varepsilon$ powers gives
$2H(\varepsilon v)=-\frac{\varepsilon(r^{3}v_{xx}+rv_{\varphi\varphi})+{\cal O}(\varepsilon^{2})}{(r^{2}+\varepsilon 2rv+{\cal O}(\varepsilon^{2}))^{3/2}}+\frac{1}{(r^{2}+\varepsilon 2rv+{\cal O}(\varepsilon^{2}))^{1/2}}.$
@@
$\frac{\,{\rm d}}{\,{\rm d}\varepsilon}2H(\varepsilon v)=-\frac{(r^{3}v_{xx}+rv_{\varphi\varphi})(r^{2}+{\cal O}(\varepsilon))^{3/2}+{\cal O}(\varepsilon)}{(r^{2}+\varepsilon 2rv+{\cal O}(\varepsilon^{2}))^{3}}-\frac{rv+{\cal O}(\varepsilon)}{(r^{2}+{\cal O}(\varepsilon))^{3/2}}.$
@@
$\frac{\,{\rm d}}{\,{\rm d}\varepsilon}K(\varepsilon v)=\frac{\,{\rm d}}{\,{\rm d}\varepsilon}\frac{-\varepsilon r^{3}v_{xx}+{\cal O}(\varepsilon^{2})}{(r^{2}+2\varepsilon 2rv+{\cal O}(\varepsilon^{2}))^{2}}=\frac{-r^{3}v_{xx}((r^{2}+\varepsilon 2rv+O(\varepsilon^{2}))^{2}+{\cal O}(\varepsilon)}{(r^{2}+\varepsilon 2rv+{\cal O}(\varepsilon^{2}))^{4}}.$
@@
$\frac{\,{\rm d}}{\,{\rm d}\varepsilon}\Delta_{\varepsilon v}H(\varepsilon v)|_{\varepsilon=0}=\left(\frac{\,{\rm d}}{\,{\rm d}\varepsilon}\Delta_{\varepsilon v}H(0)\right)\bigg{|}_{\varepsilon=0}+\Delta_{0}\left(\frac{\,{\rm d}}{\,{\rm d}\varepsilon}H(\varepsilon v)\bigg{|}_{\varepsilon=0}\right)=\Delta_{0}\left(\frac{\,{\rm d}}{\,{\rm d}\varepsilon}H(\varepsilon v)\bigg{|}_{\varepsilon=0}\right),$
since $H(0)$ is a constant and $\Delta_{\varepsilon v}$ is non degenerate for small perturbations. ∎
Inserting the expressions from Theorem 3.1 we obtain
Lemma 3.2.
@@
Over the unbounded ${\cal C}(r)$, ${\partial}_{u}G(0,\Lambda))$ has a continuous spectrum, but by restricting to the compact cylinder ${\cal C}_{L}(r)$ of length $L$, ${\partial}_{u}G(0,\Lambda))$ has a discrete spectrum.
Lemma 3.3.
The eigenvalue problem
@@
$\psi_{mn}(x,\varphi)=e^{{\rm i}(k_{m}x+n\varphi)}$, $k_{m}=2m\pi/L$, and $\mu_{mn}=\mu(k_{m},n)$ with (38a)
$\mu(k,n,\Lambda)=-\frac{1}{2}\Big{[}k^{2}{+}\frac{n^{2}}{r^{2}}\Big{]}^{2}+\left(r\lambda_{2}{+}\frac{1}{r^{2}}\right)\Big{[}k^{2}{+}\frac{n^{2}}{r^{2}}\Big{]}-\left(\frac{\lambda_{2}}{r}{+}\frac{1}{2r^{4}}\right)-\left(\frac{1}{r}{-}2c_{0}\right)\frac{k^{2}}{r}.$ (38b)
@@
we obtain (38b). Clearly $\mu(0,\pm 1,\Lambda)=0$, hence $\sin(\varphi),\cos(\varphi)\in\ker({\partial}_{u}G(0,\Lambda))$. ∎
@@
The domain of ${\partial}_{U}{\cal G}(0,\Lambda)$ is $H^{4}_{\text{per}}({\cal C}_{L}(r))\times{\mathbb{R}}^{2}$, equipped with the norm $\|U\|_{{\partial}_{(u,\Lambda)}\mathcal{G}}=\|u\|_{H^{4}}+\|\Sigma\|_{{\mathbb{R}}^{2}}$, and altogether we have the generalized eigenvalue problem
${\partial}_{U}\mathcal{G}(0,\Lambda)\Psi=\mu\begin{pmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix}\Psi,\quad\Psi=(\psi,\sigma_{1},\sigma_{2}),$ (40)
@@
$\Psi_{mn}=\begin{pmatrix}\psi_{mn}\\ 1\\ -2\end{pmatrix},\,(m,n)\in{\mathbb{Z}}^{2}\backslash\{(0,0)\},\text{ and }\Psi_{00}=\begin{pmatrix}0\\ 1\\ -2\end{pmatrix}$.
Remark 3.4.
(a) The eigenvalues $\mu(0,\pm 1,\lambda)=0$ correspond to translations of ${\cal C}_{L}(r)$ in $y$ and $z$ directions. Each eigenvalue of the form $\mu(k_{m},0,\Lambda)$ and $\mu(0,n,\Lambda)$ has multiplicity two and every “double” eigenvalue $\mu(k_{m},n,\Lambda)$ with $m\neq 0\neq n$ has multiplicity four with eigenfunctions $\psi_{\pm m\pm n}$.
(b) The eigenvalues $\mu(k_{m},n,\Lambda)$ do not explicitly depend on $\lambda_{1}$, but implicitly. Recall that $\lambda_{1}=\frac{1}{2}\left(\frac{1}{2r^{2}}-2r\lambda_{2}-2c_{0}^{2}\right)$, hence the parameters are not independent on the cylinder.